OpenStudy (anonymous):

Help! Stats! A sample size of 7 (n=7) has a mean of M=9. One of the sample scores is changed from x=19 to x=5. What is the value for the new sample mean? How can I do this without ALL of the variables?

OpenStudy (anonymous):

what. how can the mean possibly increase if one of the values decreases? For example, if the values were 1,2,4,8,14,15,19 (mean 9) and you change the 19 to 5, the new mean would be 7. If a value is decreased, then the mean will always decrease

OpenStudy (campbell_st):

well look at it this way Mean = sum of the score/number of scores so Sum of the score = mean x number of scores Sum = 9 x 7 so the sum of the scores is 63 now if 19 is removed subtract it 63 - 19 = 44 if 5 is added in place of 19 the sum of the scores is 44 + 5 = 49 so the new mean is Mean = 49/7 just calculate it.

OpenStudy (anonymous):

i was just wondering cause in your previous post you got a higher mean than the original after one of the sample numbers was decreased